### v1

@@ Abstract
We assess the vulnerability of multimodal large language models to misleading visualizations - charts that distort the underlying data using techniques such as truncated or inverted axes, leading readers to draw inaccurate conclusions that may support misinformation or conspiracy theories. Our analysis shows that these distortions severely harm multimodal large language models, reducing their question-answering accuracy to the level of the random baseline. To mitigate this vulnerability, we introduce six inference-time methods to improve performance of MLLMs on misleading visualizations while preserving their accuracy on non-misleading ones. The most effective approach involves (1) extracting the underlying data table and (2) using a text-only large language model to answer questions based on the table. This method improves performance on misleading visualizations by 15.4 to 19.6 percentage points. [^1] Code and data available at: github.com/UKPLab/arxiv2025-misleading-visualizations
@@
large language models, question answering, visualization
@@
Figure 1: Four examples of real-world misleading visualizations [1] with QA pairs. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
Visualizations are widely used to convey data insights efficiently. However, design flaws - whether intentional or not - can distort the correct interpretation of the underlying data [2, 1, 3, 4, 5, 6, 7]. These design flaws, or misleaders, include truncated, inverted, and dual axes, 3D effects, or inconsistent tick intervals. Misleading visualizations pose a serious threat to our society, as they are used to support misinformation and conspiracy theories [3]. Figure 1 shows real-world examples of misleading visualizations used to misinform readers on sensitive topics such as access to safe drinking water, politics, COVID-19, or abortion statistics. The CALVI test [8] demonstrated the impact of misleading visualizations in decreasing human readers’ ability to interpret the underlying data accurately.
The visualization reasoning abilities of multimodal large language models (MLLMs) have advanced rapidly, as demonstrated on the reference benchmark ChartQA [9], suggesting their potential as chart reasoning assistants. However, it remains uncertain whether they are sensitive to misleaders, as humans are. If vulnerable, MLLMs used as chart reasoning assistants risk amplifying both the spread of misinformation and human belief in it. Recent work has reported a sensitivity of GPT4 [10] to misleaders, but their analysis was limited to a single MLLM evaluated on six misleading visualizations [11].
In this study, we compare the question-answering (QA) performance of 16 MLLMs of varying sizes across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances and featuring 17 distinct misleaders, defined in Table 1. It combines the existing synthetic CALVI ($n$ = 45) [8] and CHARTOM ($n$ = 56) [12] datasets with 42 real-world examples [1] manually annotated with QA pairs; (b) a non-misleading visualization dataset ($n$ = 124), drawn from CALVI ($n$ = 15) and CHARTOM ($n$ = 56) and the real-world VLAT dataset ($n$ = 53) [13]; and (c) ChartQA, the reference non-misleading real-world benchmark for assessing MLLMs ($n$ = 2500).
@@
Our evaluation reveals that misleaders substantially degrade MLLM QA performance, reducing accuracy by up to 34.8 percentage points compared to non-misleading visualizations. To mitigate this vulnerability, we explore six inference-time correction methods that modify the input image and/or prompt without requiring additional training data or fine-tuning. These methods aim to reduce the impact of misleaders while minimizing any negative effect on accuracy for non-misleading visualizations. We compare the following methods, illustrated in Figure 2: (1) including a warning message in the prompt to alert the MLLM to specific misleaders in the visualization; (2) extracting the axes using the MLLM and incorporating them into the prompt; (3) extracting the underlying data table and incorporating it into the prompt; (4) combining both extracted axes and table into the prompt; (5) providing the extracted table without the visualization to a text-only LLM, reframing the task as table-based QA; and (6) using a text-only LLM to generate code for a new visualization based on the extracted table, replacing the original one as input.
@@ 1 Results
1.1 Assessing the vulnerability to misleading visualizations
Figure 3 presents the QA accuracy of 16 MLLMs across the three datasets. The results reveal three key trends.
First, MLLMs perform substantially worse on misleading visualizations than on non-misleading ones, with accuracy dropping by up to 34.8 percentage points. The decline is even more pronounced compared to ChartQA, reaching up to 65.5 percentage points. Furthermore, the mean MLLM accuracy on misleading visualizations (24.8%) is close to the random baseline (25.6%). This suggests that MLLMs struggle to correctly interpret the distorted data, making them highly vulnerable to misleaders. This aligns with findings from the CALVI study [8], where human accuracy averaged 80% on non-misleading visualizations but dropped to 39% on misleading ones.
@@ 1.1 Assessing the vulnerability to misleading visualizations
Figure 3: Top: Accuracy (%) of various MLLMs on misleading visualization [8, 12, 1], non-misleading visualization [8, 12, 13], and ChartQA datasets [9]. The horizontal dashed line indicates the accuracy of the random baseline on misleading visualizations. Models are sorted by increasing accuracy on ChartQA. Bottom: Accuracy (%) of various MLLMs on subsets of the misleading visualizations.
Second, accuracy on misleading visualizations does not follow the upward performance trend observed in ChartQA, and to a lower extent, non-misleading visualizations. From these results, we conclude that reducing MLLM vulnerability to misleading visualizations will not happen naturally as a by-product of improving performance on standard benchmarks like ChartQA. This makes dedicated mitigation methods all the more necessary.
Third, the best-performing MLLMs on misleading visualizations, GPT-4o and InternVL2.5-38B, primarily outperform other MLLMs on the real-world subset. We assume this is due to their large parametric knowledge, which spans beyond the end date of the real-world subset (2022). This knowledge about real-world events and statistics allows them to answer some questions without leveraging the visualization. However, this advantage does not extend to synthetic data in CALVI and CHARTOM.
1.2 Mitigating the impact of misleading visualizations
Figure 4: Change ($\Delta$) in accuracy (percentage points) on misleading and non-misleading visualization datasets using different inference-time correction methods to mitigate misleading visualizations. Statistically significant changes ($\leq$0.05) are hashed.
@@ 1.2 Mitigating the impact of misleading visualizations
We evaluate the impact of the correction methods with three open-weight, mid-sized MLLMs which all perform below the random baseline on misleading visualizations: Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B. Figure 4 presents the change in accuracy across the six correction methods: (1) inclusion of a warning message, inclusion of the (2) axes, (3) table, or (4) both, (5) table-based QA, and (6) redrawing the visualization. Qwen2.5-7B [14] serves as the text-only LLM for table-based QA and redrawing the visualization.
The most effective approach by far is table-based QA, yielding significant improvements of 15.4 to 19.6 percentage points. While incorrect or incomplete table extractions decrease performance on non-misleading visualizations — particularly for line charts, scatter plots, and maps — these losses are not significant.
Another promising correction method is redrawing the visualization. It achieves significant but more modest improvements for two MLLMs. This approach is effective only when the generated code compiles on a Python interpreter; otherwise, it defaults to using the original visualization. Across both datasets, the lowest redrawing success rates are observed for scatter plots (79%), and stacked bar charts (80%), which we explain by their high number of visual items and complex layouts, while the success rates for all other chart types are above 90%.
Other correction methods do not yield significant improvements. Including the extracted table in the prompt even has a significant negative impact on non-misleading visualizations for one MLLM.

### v3

@@ Abstract
Visualizations play a pivotal role in daily communication in an increasingly data-driven world. Research on multimodal large language models (MLLMs) for automated chart understanding has accelerated massively, with steady improvements on standard benchmarks. However, for MLLMs to be reliable, they must be robust to misleading visualizations, charts that distort the underlying data, leading readers to draw inaccurate conclusions that may support disinformation. Here, we uncover an important vulnerability: MLLM question-answering accuracy on misleading visualizations drops on average to the level of a random baseline. To address this, we introduce the first inference-time methods to improve performance on misleading visualizations, without compromising accuracy on non-misleading ones. The most effective method extracts the underlying data table and uses a text-only LLM to answer the question based on the table. Our findings expose a critical blind spot in current research and establish benchmark results to guide future efforts in reliable MLLMs. [^1] Code and data available at: github.com/UKPLab/arxiv2025-misleading-visualizations
@@
large language models, chart understanding, visualization
In an increasingly data-driven world, visualizations are widely used by scientists, journalists, governments, or companies to efficiently communicate data insights to a broad audience [1]. They play a crucial role in crisis settings, such as during the COVID-19 pandemic, where charts were shared daily to inform the public about the spread of the disease [2]. In many cases, visualizations support a message more convincingly than if the underlying data table was shown directly to readers [3]. However, visualizations can also be misleading [4]. This occurs when design flaws, or misleaders, distort the underlying data in a way that prevents a correct interpretation by the reader [5, 6, 4, 7, 8, 9, 10, 11, 12]. These misleaders appear across a wide range of visualizations, including bar, pie, and line charts, choropleth maps, or scatter plots. Recent taxonomies have documented over 70 distinct types of misleaders observed in real-world examples [7, 9]. Major categories include axes manipulations, such as truncated and inverted axes, or the use of visual manipulations like area encoding and 3D effects. Figure 1 shows three real-world examples of misleading visualizations [7] paired with multiple-choice questions. Given a question about the underlying data table, their misleaders lead readers to infer a misleading answer (in purple) instead of the correct one (in green). Figure 2 illustrates how two visualizations of the same data, one non-misleading and the other misleading, can lead to different interpretations by readers.
@@
Figure 1: Three examples of real-world misleading visualizations [7] with multiple choice questions. The correct answer is colored in green, while the wrong answer supported by the misleader is colored in purple.
Figure 2: A pair of bar charts, one non-misleading and one misleading, representing the same underlying data [11]. Both are paired with the same Likert-scale question, where 1 means “a little” and 6 means “a lot”. A consistent interpretation requires identical responses for both visualizations. Assuming a score of 3 is given for the non-misleading chart, only a score of 3 is valid for the misleading one. However, the truncated axis deceives readers in choosing a value higher than 3.
Prior work has revealed the potential of misleading visualizations to deceive human readers. Some studies [6, 13, 14, 15] have shown that such visualizations reduce human accuracy when answering multiple-choice, free-text, or rank questions about the underlying data. Other studies [6, 16, 11] have demonstrated that human readers tend to provide different Likert-scale answers to a question depending on whether they view a misleading or non-misleading visualization of the same data table.
Misleading visualizations pose a serious threat to society. Due to their high deceptive potential, they can be exploited by malicious actors to promote online disinformation [4]. For instance, the visualizations in Figure 1 mislead readers on sensitive topics such as COVID-19, abortion, and gun violence statistics. Misleading visualizations can shift public opinion, even on highly polarized political debates like Brexit [17].
Multimodal large language models (MLLMs) have sparked a massive interest in automated chart understanding research [1], with steady progress on reference benchmarks such as ChartQA [18, 19], highlighting their strong potential as interactive assistants for extracting data insights. However, existing research has largely overlooked the important threat posed by misleading visualizations. If MLLMs, like humans, are easily deceived by misleading visualizations, malicious actors could exploit this vulnerability. MLLMs confronted to misleading visualizations could present incorrect interpretations of the underlying data to users, contributing to the spread of disinformation, and reinforcing human belief in it. These risks underscore the urgency of thoroughly assessing the vulnerability of MLLMs to misleading visualizations and developing effective mitigation strategies.
In this work, we conduct the first comprehensive study to assess and mitigate the vulnerability of 18 MLLMs of varying sizes to misleading visualizations. This collection includes 11 open-weight MLLMs from the Llava-Next [20], Qwen2VL [21], Ovis-1.6 [22], and InternVL2.5 [23] families, ranging from 2 to 38 billion parameters. Additionally, we include four commercial models, GPT4 [24], GPT4o [24], Gemini-1.5-flash [25], and Gemini-1.5-pro [25], as well as three open-weight MLLMs specifically trained for visualization understanding: Llava-Chart-Instruct [26], TinyChart [27], and ChartGemma [28] ranging from 3 to 13 billion parameters.
In Experiment 1, we compare the question-answering (QA) accuracy of MLLMs across three datasets: (a) a misleading visualization dataset containing $n$ = 143 instances, and featuring 17 types of misleaders, defined in Table 1 [13, 14, 15, 7]. It combines the existing synthetic CALVI ($n$ = 45) [13] and CHARTOM ($n$ = 56) [14, 15] datasets with 42 real-world examples [7] that we manually annotated with multiple choice questions; (b) a non-misleading visualization dataset ($n$ = 124) [13, 14, 15, 29], also combining real-world and synthetic cases; and (c) ChartQA [18], the reference real-world non-misleading benchmark ($n$ = 2500). In Experiment 2, we evaluate the consistency of MLLMs, i.e., whether they provide the same response to a Likert-scale question when shown either a misleading or a non-misleading visualization of the same underlying data. We use the synthetic dataset of Lauer & O’Brien, which consists of four such visualization pairs [11]. Detailed statistics about the datasets, including the number of misleader categories and the type of questions, are reported in Table 2.
Experiments 1 and 2 demonstrate that MLLMs, like humans, are indeed vulnerable to misleading visualizations. Experiment 1 yields several key insights: QA accuracy drops by up to 65.5 percentage points (pp) compared to ChartQA, with half the MLLMs failing to outperform the random baseline; MLLMs are less vulnerable to real-world misleading visualizations than to synthetic ones; and vulnerability varies substantially depending on the type of misleader.In addition, Experiment 2 shows that MLLMs exhibit inconsistency when assigning Likert-scale scores to two visualizations of the same data when one of them has a truncated axis.
To mitigate the high vulnerability of MLLMs observed in Experiment 1, we introduce six inference-time correction methods in Experiment 3. These methods modify the input image and/or prompt without requiring additional training data or fine-tuning. We evaluate them on both the misleading and non-misleading visualization datasets used in Experiment 1. The goal is to improve QA accuracy on misleading visualizations while maintaining performance on non-misleading ones.
We compare the following methods, illustrated in Figure 3: (1) including a warning message to the prompt to make the MLLM aware of specific misleaders in the visualization; (2) extracting the axes using the MLLM and incorporating them into the prompt; (3) extracting the underlying data table and adding it into the prompt; (4) combining both extracted axes and table into the prompt; (5) providing the extracted table, without the visualization, to a text-only LLM, reframing the task as table-based QA; and (6) using a text-only LLM to generate code for a new visualization based on the extracted table, which then replaces the original visualization in the QA prompt.
In summary, our contributions are twofold: (1) We present the first extensive analysis of the vulnerabilities of 18 MLLMs to 17 types of misleading visualizations, evaluating both QA accuracy and Likert-scale consistency. (2) We are the first to introduce six correction methods to mitigate the negative impact of misleaders on QA performance, while maintaining accuracy on non-misleading visualizations.
The results of the three experiments expose a critical blind spot in current automated chart understanding research, highlighting the need for future work to make MLLMs more robust to misleading visualizations.
@@ 1 Results
1.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Experiment 1 assesses the ability of MLLMs to answer questions about the underlying data given a visualization. The upper part of Figure 4 presents the QA accuracy of 18 MLLMs across the three datasets, sorted by increasing accuracy on ChartQA. The bottom part displays the accuracy for the three subsets constituting the misleading visualizations dataset. The results of Experiment 1 reveal three key findings.
@@ 1.1 Experiment 1 - Assessing MLLM vulnerabilities (Accuracy)
Figure 4: Top: Accuracy (%) of various MLLMs on misleading visualization [13, 14, 15, 7], non-misleading visualization [13, 14, 15, 29], and ChartQA datasets [18]. The horizontal dashed line indicates the accuracy of the random baseline on misleading visualizations. Models are sorted by increasing accuracy on ChartQA. Bottom: Accuracy (%) of various MLLMs on subsets of the misleading visualizations. The horizontal dashed lines indicate average human accuracy on CALVI and CHARTOM [13, 14, 15].
1.1.1 Performance is worse on misleading visualizations
All MLLMs perform substantially worse on misleading visualization compared to the non-misleading visualizations dataset, with accuracy dropping by up to 34.8 pp, with an average of 23.4 pp. The decline is even more pronounced when compared to ChartQA, reaching up to 65.5 pp, with an average of 54.0 pp. Furthermore, the mean MLLM accuracy on misleading visualizations (25.6%) is on par with the random baseline score (25.6%). These results demonstrate that MLLMs struggle to correctly interpret the distorted data, making them highly vulnerable to misleaders.
For comparison, average human accuracy reported in prior work on CALVI and CHARTOM is 80 and 89% on non-misleading visualizations, dropping to 39 and 49% on misleading ones, respectively [13, 14, 15]. On the same datasets, MLLMs achieve an average of 51 and 47% on non-misleading visualizations and 24 and 21% on misleading ones, falling short of human performance in all cases.
The ranking of MLLMs by accuracy varies a lot across the three datasets. Notably, the two best-performing models on the CALVI subset of misleading visualizations rank among the worst on ChartQA, placing second and third from the bottom. This discrepancy suggests that achieving high accuracy on misleading visualizations is unlikely to emerge naturally from improvements on ChartQA, underscoring the need for dedicated mitigation methods.
1.1.2 Performance is higher on real-world misleading visualizations
MLLMs perform better on the real-world subset of the misleading visualizations dataset, achieving an average accuracy of 33.5%, compared to 24% on CALVI and 21% on CHARTOM. In general, the best-performing MLLMs on misleading visualizations, GPT4o, Gemini-1.5-pro, and InternVL2.5-38B, primarily distinguish themselves through their higher performance on this subset. We hypothesize that this advantage stems from their parametric knowledge, which extends beyond 2022, the endpoint of the real-world subset. This allows these models to answer some questions based on stored knowledge about past events and statistics, without relying on the content of the visualization. In such cases, parametric knowledge effectively serves as a form of protection against real-world misleaders. However, this finding should be interpreted with caution. In practical scenarios, the underlying data will often be recent and therefore unlikely to be covered by the model’s parametric knowledge.
1.2 Performance varies by type of misleader
Figure 5: Average Likert-scale ratings (1 to 6) of MLLMs and human readers for the four pairs of misleading and non-misleading visualizations depicting the same underlying data from the dataset of Lauer & O’Brien [11]. MLLM results are reported with standard deviations. Human results are reported from prior work [11].
Not all misleading visualizations are equally deceptive. This was already observed with human readers [13, 14, 15], and is here further confirmed with MLLMs. Detailed accuracy scores per misleader and MLLM are reported in Figure 7. The three misleaders on which MLLMs perform best are concealed uncertainty, misleading annotations, and misrepresentation, with average accuracies of 64.8, 50.0, and 43.7, respectively. The four on which they perform worse are overplotting and inappropriate use of pie charts, both at 0.0%, and area encoding and missing data at 5.6%. For the two most represented misleaders, truncated axis and inverted axis, the MLLMs achieve average accuracies of 23.5 and 27.4 %, respectively.
There are notable differences in which misleaders pose greater challenges to MLLMs compared to humans, and vice versa. For instance, area encoding is highly misleading for MLLMs, but not for human readers, who achieve an average accuracy of 91.5% on this type of misleader [14]. Conversely, both humans and MLLMs struggle with misleaders such as overplotting, missing data, and 3D effects [13]. The use of a dual axis is a misleader that poses greater difficulty for humans, who score an average accuracy of 16.1% on CHARTOM [14], compared to MLLMs, which achieve an average of 35.4%. These results highlight a notable difference in how humans and MLLMs process visualizations and are deceived by misleaders. This suggests that the methods needed to make MLLMs more robust to misleading visualizations will, at least in part, differ from those designed to protect human readers [30].
1.3 Experiment 2 - Assessing MLLM vulnerabilities (Consistency)
Experiment 2 evaluates whether MLLMs interpret consistently two visualizations that represent the same underlying data table, one misleading and the other non-misleading. This is assessed using a Likert-scale question, such as in Figure 2. Unlike in Experiment 1, there is no ground truth answer. Instead, since both visualizations depict the same underlying data, a model is considered misled if it assigns different scores to each of them. We assess the significance of the difference in ratings with a two-sided paired t-test (p-value (p) $\leq$ 0.05).
Figure 5 summarizes the average Likert-scale ratings from all 18 MLLMs for each visualization pair. Human results from prior work [11] are included for comparison. The average rating difference between the misleading and non-misleading visualizations is similar for humans and MLLMs, though MLLMs tend to assign slightly higher scores overall. For MLLMs, the inconsistency in ratings is statistically significant for the two visualization pairs involving truncated axis: one with bar charts (Test statistic (T) = -3.2, p = 5e-3) and one with line charts (T = -3.0, p = 0.01). For the other two pairs, misrepresentation with bubble charts (T = -0.77, p = 0.45) and 3D effects with pie charts (T = -0.62, p = 0.55), the differences are not statistically significant. These results further support the conclusion that MLLMs are vulnerable to misleading visualizations, raising questions about their ability to consistently recognize the same underlying data table behind a misleading and a non-misleading version, in particular for truncated axis. However, this vulnerability must be interpreted with nuance, as not all misleaders showed a significant impact on MLLM’s answers consistency.
1.4 Experiment 3 - Correction methods
@@
Having established that MLLMs are vulnerable to misleading visualizations, we explore six correction methods aimed at improving their robustness. Our focus is on increasing accuracy on the misleading visualizations dataset from Experiment 1, while maintaining performance on the non-misleading dataset. We evaluate the impact of six correction methods using three open-weight, mid-sized MLLMs, Qwen2VL-7B, Ovis1.6-9B, and InternVL2.5-8B, all of which performed below the random baseline on misleading visualizations in Experiment 1. Statistical significance of the changes in accuracy is assessed using the two-sided McNemar test (p $\leq$ 0.05) [31]. All test statistics and p-values are reported in Table 3. Figure 6 shows the change in accuracy relative to Experiment 1 across the six methods on the misleading and non-misleading visualization datasets: (1) inclusion of a warning message, (2) inclusion of extracted axes, (3) inclusion of the extracted table, (4) inclusion of both axes and table, (5) table-based QA, and (6) redrawing the visualization. Qwen2.5-7B [32] is used as the text-only LLM for both table-based QA and redrawing visualizations.
Figure 6: Change ($\Delta$) in accuracy (percentage points) compared to Experiment 1 on misleading and non-misleading visualization datasets using different inference-time correction methods to mitigate misleading visualizations. Statistically significant changes ($\leq$0.05) are hashed.
1.4.1 Two methods are effective: table-based QA and redrawing the visualization
The most effective approach by far to counter misleaders is table-based QA, which yields significant improvements on misleading visualizations, ranging from 15.4 to 19.6 pp. Its strength lies in fully replacing the visualization with a data table, thereby eliminating misleaders that do not transfer to the tabular format, such as inverted and truncated axis, or misleading annotations. However, this method proves less effective against misleaders that persist even when the data is presented in table form, such as inappropriate item order and concealed uncertainty. Table-based QA comes with a notable cost for non-misleading visualizations. Two models experience significant decreases in accuracy, by up to 8.5 pp. This degradation is primarily due to errors in the intermediate table extraction step, where incorrect values are introduced or relevant information is lost.
Another promising correction method is redrawing the visualization. This method aligns with prior efforts to make human readers more robust to misleading visualizations [30]. It leads to significant, though more modest, improvements for two MLLMs. The visualization is redrawn using the default settings of the Python library Matplotlib [33], which inherently avoids certain misleaders like inverting the axes or using inconsistent tick intervals. In Figure 3, the redrawn visualization corrects the misleading trend line by eliminating inconsistent tick intervals, leading to a more accurate representation of the data. This correction method is effective only if the generated code compiles successfully on a Python interpreter; otherwise, the original visualization is used. The lowest redrawing success rates are observed for scatter plots (79%), and stacked bar charts (80%), likely due to their higher visual complexity and number of elements, which require Qwen2.5-7B to produce longer and more detailed Python code. For all other chart types, the redrawing success rate exceeds 90%. Unlike table-based QA, this method does not result in significant decreases in accuracy on non-misleading visualizations, making it a more attractive option than table-based QA when preserving performance on non-misleading visualization is a priority.
1.4.2 Table extraction is a key intermediate step
Both table-based QA and redrawing the visualization depend heavily on the accuracy of an intermediate step: table extraction. Since the underlying data table is not available, it must be derived from the visualization before either correction method can be applied. This extraction step is often non-trivial, particularly for complex chart types such as maps or scatter plots that contain a high number of visual elements to extract. Additionally, misleaders themselves can degrade MLLMs’ ability to extract data accurately. Perfect accuracy in table extraction is not always necessary, its importance depends on the nature of the question. Some questions can be answered correctly as long as the extracted table preserves the overall trend of the underlying data, while others require exact values. As discussed above, errors in table extraction can significantly decrease accuracy on non-misleading visualizations. In the absence of ground truth tables, we cannot directly evaluate the accuracy of table extraction. Instead, we analyze MLLM outputs on CHARTOM [14, 15], which pairs each question with two visualizations, one misleading and one non-misleading, generated from the same underlying data. Ideally, the tables extracted from both visualizations should be identical. However, for all MLLMs, the extracted tables have a perfect match in only 4 out of 56 pairs (7.1%), and a partial match for 13 to 14 other pairs (23 to 25%). Figure 8 illustrates the extraction of two different tables from a CHARTOM pair. Given that the two most effective correction methods depend on reliable table extraction, and that extraction errors also negatively impact performance on non-misleading visualizations, improving this intermediate step is a key direction for future work.
1.4.3 Other correction methods are not effective